OpenStudy (anonymous):

Find the geometric mean of 7 and 9. A. 2004-04-01-01-00_files/i0220000.jpg B. 2004-04-01-01-00_files/i0220003.jpg C. 2004-04-01-01-00_files/i0220001.jpg D. 2004-04-01-01-00_files/i0220002.jpg

OpenStudy (anonymous):

um.... What are the answer choices?

OpenStudy (anonymous):

a. 3 square root of 7 b. square root of 70 c.6 square root of 2 d. 63

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So if you do the math you end up with \[\sqrt{63}\]

OpenStudy (anonymous):

And do you know what the squareroot of 63 is??

OpenStudy (anonymous):

And no its not D

OpenStudy (anonymous):

7.9

OpenStudy (anonymous):

(not sure of the actuall name) but you have to tree factor since you cant end up with a decimal

OpenStudy (anonymous):

geometric mean of n numbers is given by \[G_{n}=(a_{1} \times a_{2} \times a_{3} \times ..... \times a_{n})^{\frac{1}{n}}\] for 2 numbers, \[G_{2}=(a_{1} \times a_{2})^{\frac{1}{2}}=\sqrt{a_{1}a_{2}}\]

OpenStudy (anonymous):

im still lost